1.If 100 people are selected at random, what is the chance that not more than ten people suffer from COVID 19-disease

2. Does talking on mobile phones while driving distract people? Researchers measured the reaction times of 80 study

participants as they talked on mobile phones and found that the average level of distraction from their driving was

rated to be 60 percent. Explain what could be your population, Sample, sample statistic and characteristic.


1
Expert's answer
2021-05-04T11:27:40-0400

1.) We have selected 100 random people,

Probability that the person suffer from Covid disease = "\dfrac{1}{2}"

We have to find,

The probability that not more than ten people suffer from COVID 19-disease

"1-P(X>10)"


"= P(X\le 10)"


"= P(X=0)+P(X=1)+P(X=2)+P(X=3)+P(X=4)+P(X=5)+P(X=6)+P(X=7)+P(X=8)+P(X=9)+P(X=10)"


"= (\dfrac{1}{2})^{10} [\binom{10}0+\binom{10}1+\binom{10}2+\binom{10}3+\binom{10}4+\binom{10}5+\binom{10}6+\binom{10}7+\binom{10}8+\binom{10}9+\binom{10}{10}]"


2.) The population could be people that talk on mobile phones while driving.

The sample is 80 people of the study.

The sample statistic is the average level of distraction from their driving - the mean is 60%.